# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ Abstract.
We investigate the nearsightedness property in the linear tight binding model at zero Fermi-temperature. We focus on the decay property of the density matrix for materials with indirect band gaps. By representing the density matrix in reciprocal space, we establish a qualitatively sharp estimate for the exponential decay rate in homogeneous systems. An extending result under perturbations is also derived. This work refines the estimates presented in (Ortner, Thomas & Chen 2020), particularly for systems with small band gaps.
@@ 1. Introduction
The electronic structure determines a wide range of physical and chemical properties of materials. For its computation, density functional theory (DFT) has been successfully applied for many decades [13, 23]. However, even in the minimalistic tight binding models, the practical implementations typically incur a high computational cost, scaling cubically with the system size. One approach to alleviate this high cost is through linear scaling algorithms [2, 15, 22, 25], which rely on a notion of locality of electronic systems, often known as “nearsightedness”.
@@
Therefore, exploring the dependence of the decay rate on various factors is important and has provoked extensive discussion, both analytically and numerically [1, 5, 14, 18, 27]: it is determined by the band gap in insulators and by the Fermi-temperature in metals. However, these results may not always be sufficiently precise to account for the exponential decay observed in some materials with small band gaps, such as semiconductors like silicon and germanium. Numerical results from [27] show that in the silicon system, the decay is nearly identical to the carbon system while the former has a much smaller band gap.
The main purpose of this paper is to explore a subtle point that, to the best of our knowledge, has been missed in the nearsightedness discussion to date: the decay rate of the density matrix depends not on the direct band gap (the most common notion of a band gap) but on the indirect band gap; cf. Figures 1 and 2.
@@
This result is important because the indirect band gap can be much larger than the direct band gap, especially in small-gap semiconductors such as ${\rm Mg}_{2}{\rm Si}$; cf. Figure 2. The result explains why some small gapped systems still exhibit excellent exponential localization and linear scaling methods remain applicable.
Next, we explore whether such improved locality estimates also generalize to the interatomic forces, which is called strong locality in [5, 26]. Through a combination of analytical and numerical empirical results we show that this is not the case. This negative result demonstrates that locality of interatomic forces is a fundamentally stronger concept than classical nearsightedness.
@@ 1.1. Outline
In Section 2, we introduce the physical background and linear tight binding models. We then provide definitions for the direct band gap $\mathsf{gap}_{-}$ and indirect band gap $\mathsf{gap}_{+}$ in reciprocal space. In Section 3, we state the main results of this paper. To this end, we first discuss the dependence of nearsightedness in homogeneous systems, and then extend the result to the inhomogeneous systems with bounded perturbations. The results show that the decay rate actually depends on $\mathsf{gap}_{+}$, and perturbations introduce a slower decay controlled by $\mathsf{gap}_{-}$ and the norm of the perturbations. At the end of our analysis, we further explore the decay rate for the derivatives of the density matrix, showing it depends on $\mathsf{gap}_{-}$. In Section 4, we provide numerical experiments to verify our analytical results. The main conclusions of this paper are summarised in Section 5. All proofs are collected in Section 6. In the appendix, we provide details of the Bloch transform, along with the specifics of numerical experiments.
@@ 1.2. Notation
[…]
@@ 2.1. Atomistic Model and Admissible Configurations
$\Lambda:=\bigcup_{i=1}^{M}\big{(}p_{i}+\mathsf{A}\mathbb{Z}^{d}\big{)}$ (2.1)
@@
$\bm{r}_{\ell k}(u):=\ell+u(\ell)-k-u(k)\quad\text{and}\quad r_{\ell k}(u):=|\bm{r}_{\ell k}(u)|.$ (2.2)
@@
$D_{\rho}u(\ell)\coloneqq u(\ell+\rho)-u(\ell),\quad\forall\ell\in\Lambda,~{}\rho\in\Lambda-\ell.$ (2.3)
@@
$\|Du\|_{\ell_{\Upsilon}^{2}}:=\left(\sum_{\ell\in\Lambda}\left|Du(\ell)\right|^{2}_{\Upsilon}\right)^{1/2}~{}~{}\text{with}~{}~{}\left|Du(\ell)\right|^{2}_{\Upsilon}:=\sum_{\rho\in\Lambda-\ell}e^{-2\Upsilon|\rho|}\left|D_{\rho}u(\ell)\right|^{2}.$ (2.4)
@@
$\mathrm{Adm}(\Lambda):=\left\{u:\Lambda\rightarrow\mathbb{R}^{d}\colon\|Du\|_{\ell_{\Upsilon}^{2}}<\infty,\,\,\exists\,\mathfrak{m}>0~{}\text{with}~{}r_{\ell k}(u)\geq\mathfrak{m}\,\,\forall\ell,k\in\Lambda\right\}.$ (2.5)
@@ 2.2. Tight Binding Model
${\mathcal{H}}(u)_{\ell k}^{ab}=h^{ab}\big{(}{\bm{x}}_{\ell k}(u)\big{)}+\delta_{\ell k}\sum_{m\neq\ell}t^{ab}({\bm{x}}_{\ell m}(u))$ (2.6)
for $\ell,k\in\Lambda$ and $1\leq a,b\leq N_{\mathrm{b}}$. The functions $h^{ab},t^{ab}:(\mathbb{R}^{d},~{}\{z_{\ell}\}_{\ell\in\Lambda},~{}\{z_{\ell}\}_{\ell\in\Lambda})\rightarrow\mathbb{R}$ are continuously differentiable with
$\left|h^{ab}({\bm{x}}_{\ell k})\right|+\left|\nabla h^{ab}({\bm{x}}_{\ell k})\right|\leq h_{0}e^{-\gamma_{0}r_{\ell k}}\quad\text{and}$
$\left|t^{ab}({\bm{x}}_{\ell k})\right|+\left|\nabla t^{ab}({\bm{x}}_{\ell k})\right|\leq h_{0}e^{-\gamma_{0}r_{\ell k}}$ (2.7)
@@
Remark 2.1. We note that the constants $h_{0},\gamma_{0}$ in (2.2) may be chosen to be independent of atomic species. Additionally, (2.2) requires that the dependence of ${\mathcal{H}}(u)^{ab}_{\ell k}$ on site $m$ decays exponentially in $r_{\ell m}(u)+r_{mk}(u)$. That is, we are assuming long-range Coulomb interactions are screened, a common assumption in many practical tight binding models [9, 24, 29]. We assume this to separate Coulomb type long-range range effects from those arising due to small band gaps.
@@ 2.3. Direct and Indirect Band Gaps
Next, we review the concepts of direct and indirect band gaps. For this purpose, we first recall the Bloch transform, which provides a decomposition of ${\mathcal{H}}^{\rm ref}$, denoted by ${\mathcal{H}}_{\boldsymbol{\xi}}$ for all $\bm{\xi}\in\Gamma^{*}$. For more details, see Appendix A. By applying an eigenvalue decomposition for ${\mathcal{H}}_{\boldsymbol{\xi}}$, we can obtain energy band structure of the system and subsequently distinguish between direct and indirect band gaps.
$\big{[}{\mathcal{H}}_{\boldsymbol{\xi}}\big{]}^{ab}_{\ell_{0}k_{0}}\coloneqq\sum_{\alpha\in\mathbb{Z}^{d}}[{\mathcal{H}}^{\mathrm{ref}}]^{ab}_{\ell_{0}+\mathsf{A}\alpha,k_{0}}e^{-i(\ell_{0}-k_{0}+\mathsf{A}\alpha)\cdot\bm{\xi}}\quad\forall\bm{\xi}\in\Gamma^{*},$ (2.8)
for all $\ell_{0},k_{0}\in\Lambda_{0}\coloneqq\Gamma\cap\Lambda,1\leq a,b\leq N_{\mathrm{b}}$. The size of the matrix ${\mathcal{H}}_{\boldsymbol{\xi}}$ is $MN_{\mathrm{b}}\times MN_{\mathrm{b}}$. For $\ell\in\Lambda$, we write $\ell_{0}$ for the unique site in $\Lambda_{0}$ satisfying $\ell-\ell_{0}\in\mathsf{A}\mathbb{Z}^{d}$. Denoting the eigenvalues of ${\mathcal{H}}_{\boldsymbol{\xi}}$ by $\varepsilon_{1}(\bm{\xi})\leq\cdots\leq\varepsilon_{MN_{\mathrm{b}}}(\bm{\xi})$, we have
$\sigma({\mathcal{H}}^{\mathrm{ref}})=\bigcup_{j=1}^{MN_{\mathrm{b}}}\left\{\varepsilon_{j}(\bm{\xi}):~{}\bm{\xi}\in\Gamma^{*}\right\},$ (2.9)
@@
We focus exclusively on systems with a positive band gap, such as insulators and semi-conductors, where the Fermi level, $\varepsilon_{\mathrm{F}}$, lies inside the band gap. The Fermi level $\varepsilon_{\mathrm{F}}$ separates the energy bands into the valence bands and conduction bands [10]; that is,
$\varepsilon_{\mathrm{F}}\not\in\sigma({\mathcal{H}}^{\mathrm{ref}})\quad\text{and}\quad\varepsilon_{N_{0}}(\bm{\xi})<\varepsilon_{\mathrm{F}}<\varepsilon_{N_{0}+1}(\bm{\xi})\quad\forall\bm{\xi}\in\Gamma^{*},$ (2.10)
@@
In semiconductors, one can consider two types of band gaps: the direct gap $\mathsf{gap}_{-}$ and the indirect gap $\mathsf{gap}_{+}$, defined respectively as
$\mathsf{gap}_{+}$ $\coloneqq\min_{\bm{\xi}\in\Gamma^{*}}\big{(}\varepsilon_{N_{0}+1}(\bm{\xi})-\varepsilon_{N_{0}}(\bm{\xi})\big{)},\qquad\text{and}$ (2.11)
$\mathsf{gap}_{-}$ $\coloneqq\min_{\bm{\xi}\in\Gamma^{*}}\varepsilon_{N_{0}+1}(\bm{\xi})-\max_{\bm{\xi}\in\Gamma^{*}}\varepsilon_{N_{0}}(\bm{\xi}).$ (2.12)
One can readily verify the following relation
$\mathsf{gap}_{+}=\min_{\bm{\xi}\in\Gamma^{*}}\big{(}\varepsilon_{N_{0}+1}(\bm{\xi})-\varepsilon_{N_{0}}(\bm{\xi})\big{)}~{}\geq~{}\min_{\bm{\xi}\in\Gamma^{*}}\varepsilon_{N_{0}+1}(\bm{\xi})-\max_{\bm{\xi}\in\Gamma^{*}}\varepsilon_{N_{0}}(\bm{\xi})=\mathsf{gap}_{-}.$ (2.13)
@@
Figure 1. Schematic of the direct and indirect band gaps.
Figure 2. The band structure of $\rm{Mg_{2}Si}$ and $\rm{C}$. In Carbon, we have $\mathsf{gap}_{-}\approx\mathsf{gap}_{+}$, whereas in $\rm{Mg_{2}Si}$, we observe $\mathsf{gap}_{-}\ll\mathsf{gap}_{+}$.
@@ 3. Main Results
3.1. Nearsightedness
We now turn the attention to the nearsightedness property, which we sometimes call weak locality to distinguish it from locality of the mechanical response (strong locality). We first focus on the homogeneous lattice, and present an improved estimation for the locality of the density matrix, explicitly tracking the dependence on $\mathsf{gap}_{+}$. Subsequently, we extend the analysis to include inhomogeneous cases with perturbations.
At zero Fermi-temperature, the Fermi-Dirac function is given by $\chi_{(-\infty,\varepsilon_{\mathrm{F}})}$ and thus the density matrix reads
$\rho(u)=\chi_{(-\infty,\varepsilon_{\mathrm{F}})}({\mathcal{H}}(u)),\quad u\in\mathrm{Adm}(\Lambda).$ (3.1)
For simplicity, we will denote by $\rho^{\mathrm{ref}}:=\rho(\bm{0})$, for which we have the following result:
Theorem 3.1. Suppose $\Lambda$ has the form of (2.1) and ${\mathcal{H}}^{\mathrm{ref}}$ satisfies (TB). Then, there exist constants $C_{1},\eta_{+}>0$ such that
$\left|\rho^{\mathrm{ref}}_{\ell k,ab}\right|\leq\frac{C_{1}}{\mathsf{gap}_{+}}e^{-{\eta_{+}}r_{\ell k}},\quad\forall\ell,k\in\Lambda,~{}1\leq a,b\leq N_{\mathrm{b}},$ (3.2)
where $\eta_{+}:=c_{1}\frac{\gamma_{0}}{h_{0}}\min\left\{h_{0},\gamma_{0}^{d}\,\mathsf{gap}_{+}\right\}$ and $c_{1},C_{1}>0$ only depend on $d,\mathfrak{m},M,N_{\mathrm{b}}$.
@@ 3.1. Nearsightedness
Remark 3.3 (Existing results). We briefly note here that (3.2) but with $\mathsf{gap}_{+}$ replaced with $\mathsf{gap}_{-}$ is classical and follows from a contour integral representation of the electron density, together with existing Combes–Thomas type estimates on the resolvent [1, 5, 12, 14, 27]. That is, we let $\mathscr{C}$ be a simple, closed, positively oriented contour encircling $\sigma({\mathcal{H}}^{\mathrm{ref}})\cap(-\infty,\varepsilon_{\mathrm{F}})$ (depicted in Figure 3), write
$\rho^{\mathrm{ref}}=\oint_{\mathscr{C}}\chi_{(-\infty,\varepsilon_{\mathrm{F}})}(z)\big{(}z-{\mathcal{H}}^{\mathrm{ref}}\big{)}^{-1}\frac{\mathrm{d}z}{2\pi i}=\oint_{\mathscr{C}}\big{(}z-{\mathcal{H}}^{\mathrm{ref}}\big{)}^{-1}\frac{\mathrm{d}z}{2\pi i},$ (3.3)
@@
$\left|\big{(}z-{\mathcal{H}}^{\mathrm{ref}}\big{)}^{-1}_{\ell k}\right|\leq\frac{2}{\mathfrak{d}^{\rm ref}}e^{-\gamma_{\mathrm{CT}}(\mathfrak{d}^{\rm ref})r_{\ell k}}$ (3.4)
where $\gamma_{\mathrm{CT}}(\mathfrak{d}^{\rm ref})=c_{0}\frac{\gamma_{0}}{h_{0}}\min\{h_{0},\gamma_{0}^{d}\,\mathfrak{d}^{\rm ref}\}$ with $c_{0}=c_{0}(d,\mathfrak{m})$ depending only on the geometry.
Figure 3. A schematic plot of the contour $\mathscr{C}$ circling the whole spectrum of valence bands denoted by the green strip, and the conduction bands denoted in blue.
Sketch of the Proof of Theorem 3.1.
@@
$\rho^{\mathrm{ref}}_{\ell k,ab}$ $=\oint_{\mathscr{C}}\bigg{[}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{\star}}\left(z-{\mathcal{H}}_{\boldsymbol{\xi}}\right)^{-1}e^{i\bm{\xi}\cdot(\ell-k)}\mathrm{d}\bm{\xi}\bigg{]}_{\ell_{0}k_{0},ab}\frac{\mathrm{d}z}{2\pi i}$
[…]
@@
where $1\leq a,b\leq N_{\mathrm{b}}$, $\ell_{0},k_{0}\in\Lambda_{0}$ satisfy $\ell-\ell_{0},k-k_{0}\in\mathsf{A}\mathbb{Z}^{d}$, and $\nu:\Gamma^{*}\rightarrow\mathbb{C}$ satisfies $\varepsilon_{N_{0}}(\bm{\xi})<\varepsilon_{\mathrm{F}}+\nu(\bm{\xi})<\varepsilon_{N_{0}+1}(\bm{\xi})$.
@@
We conclude by showing that $\nu$ and thus $\widehat{\rho}$ extend to analytic functions on a strip of width proportional to $\mathsf{gap}_{+}$ and then apply a suitable Paley–Wigner estimate. For full details, see Section 6.1. ∎
@@
We now show that the previous locality result is stable under perturbations. Specifically, we extend the result to inhomogeneous systems with finite-energy displacements of the lattice. Since those perturbations must have finite energy-norm, they are necessarily localized, but with relatively slow (e.g., algebraic) decay.

### v2

@@ Abstract.
We investigate the nearsightedness property in the linear tight binding model at zero Fermi-temperature. We focus on the decay property of the density matrix for materials with indirect band gaps. By representing the density matrix in reciprocal space, we establish a qualitatively sharp estimate for the exponential decay rate in homogeneous systems, possibly with localized perturbations. This work refines the estimates presented in (Ortner, Thomas & Chen, 2020) for systems with small band gaps.
@@ 1. Introduction
The electronic structure determines a wide range of physical and chemical properties of materials. For its computation, density functional theory (DFT) has been successfully applied for many decades [14, 24]. However, even in the minimalistic tight binding models, practical implementations typically incur a high computational cost, scaling cubically with the system size. One approach to alleviate this high cost is through linear scaling algorithms [2, 16, 23, 27], which rely on a notion of locality of electronic systems, often known as “nearsightedness”.
@@
Therefore, exploring the dependence of the decay rate on various factors is important and has provoked extensive discussion, both analytically and numerically [1, 7, 15, 19, 28]: it is determined by the band gap in insulators and by the Fermi-temperature in metals. However, these results may not always be sufficiently precise to account for the exponential decay observed in some materials with small band gaps, e.g. in semiconductors like silicon and germanium. Numerical results from [28] show that in the silicon system, the decay is nearly identical to the carbon system even though the former has a much smaller band gap.
The main purpose of this paper is to explore a subtle point that, to the best of our knowledge, has been missed in the nearsightedness discussion to date: the decay rate of the density matrix depends not on the gap (the most common notion of a band gap) but on the vertical gap; cf. Figures 1 and 2.
@@
This result is important because the vertical gap can be much larger than the gap, especially in small-gap semiconductors such as ${\rm Mg}_{2}{\rm Si}$; cf. Figure 2. The result explains why some small gapped systems still exhibit excellent exponential localization and linear scaling methods remain applicable.
Next, we explore whether such improved locality estimates also generalize to interatomic forces, which is called strong locality in [3, 29]. Through a combination of analytical and numerical empirical results we show that this is not the case. This negative result demonstrates that locality of interatomic forces is a fundamentally stronger concept than classical nearsightedness.
@@ 1.1. Outline
In Section 2, we introduce the physical background and linear tight binding models. We then provide definitions for the gap and vertical gap in reciprocal space. In Section 3, we present the main results of this work. We begin by discussing the dependence of nearsightedness principle in homogeneous systems, and then extend the analysis to strong locality by investigating the decay rate for the derivatives of the density matrix. Under additional regularity assumption, we further provide sharp estimates for both cases. Finally, we extend our analysis to inhomogeneous systems with bounded perturbations. In Section 4, we provide numerical experiments to verify our analytical results. The main conclusions of this paper are summarised in Section 5. All proofs are collected in Section 6. In the appendix, we provide details of the Bloch transform, along with the specifics of numerical experiments and estimates for the exponential sum.
@@ 1.2. Notation
The symbol $\|\cdot\|$ and $|\cdot|$ will denote the $\ell^{2}$ and Euclidean norms on $\mathbb{R}$ or $\mathbb{C}$, respectively. The subscript of $\|\cdot\|$ indicates the norm is taken in which sense. In particular, we will use $\|\cdot\|_{\rm F}$ to denote the Frobenius norm for matrices. The ball of radius $R$ about $\ell$ and $\bf 0$} will be denoted by $B_{R}(\ell)$ and $B_{R}$, respectively. For a shifted lattice, we will write $\Lambda-\ell:=\{x-\ell:x\in\Lambda\}$. For two sets $A$ and $B$, we denote the Minkowski sum of them by $A+B:=\{\bm{a}+\bm{b}~|~\bm{a}\in A,\bm{b}\in B\}$, which is formed by adding each vector in $A$ and $B$. For a self-adjoint operator $T$, the spectrum of $T$ will be denoted by $\sigma(T)$.
@@ 2.1. Atomistic Model and Admissible Configurations
(2.1) $\Lambda:=\bigcup_{i=1}^{M}\big(p_{i}+\mathsf{A}\mathbb{Z}^{d}\big)$
@@
(2.2) $\bm{r}_{\ell k}(u):=\ell+u(\ell)-k-u(k)\quad\text{and}\quad r_{\ell k}(u):=|\bm{r}_{\ell k}(u)|.$
@@
(2.3) $D_{\rho}u(\ell)\coloneqq u(\ell+\rho)-u(\ell),\quad\forall\ell\in\Lambda,~\rho\in\Lambda-\ell.$
@@
(2.4) $\|Du\|_{\ell_{\Upsilon}^{2}}:=\left(\sum_{\ell\in\Lambda}\left|Du(\ell)\right|^{2}_{\Upsilon}\right)^{1/2}~~\text{with}~~\left|Du(\ell)\right|^{2}_{\Upsilon}:=\sum_{\rho\in\Lambda-\ell}e^{-2\Upsilon|\rho|}\left|D_{\rho}u(\ell)\right|^{2}.$
@@
(2.5) $\mathrm{Adm}(\Lambda):=\left\{u:\Lambda\rightarrow\mathbb{R}^{d}\colon\|Du\|_{\ell_{\Upsilon}^{2}}<\infty,\,\,\exists\,\mathfrak{m}>0~\text{with}~r_{\ell k}(u)\geq\mathfrak{m}\,\,\forall\ell,k\in\Lambda\right\}.$
@@ 2.2. Tight Binding Model
(2.6) ${\mathcal{H}}(u)_{\ell k}^{ab}=h^{ab}\big({\bm{x}}_{\ell k}(u)\big)+\delta_{\ell k}\sum_{m\neq\ell}t^{ab}({\bm{x}}_{\ell m}(u))$
for $\ell,k\in\Lambda$ and $1\leq a,b\leq N_{\mathrm{b}}$. The functions $h^{ab},t^{ab}:\mathbb{R}^{d}\times\{z_{\ell}\}_{\ell\in\Lambda}\times\{z_{\ell}\}_{\ell\in\Lambda}\rightarrow\mathbb{R}$ are continuously differentiable (in the first variable) with
(2.7) $\left|h^{ab}({\bm{x}}_{\ell k})\right|+\left|t^{ab}({\bm{x}}_{\ell k})\right|+\left|\nabla h^{ab}({\bm{x}}_{\ell k})\right|+\left|\nabla t^{ab}({\bm{x}}_{\ell k})\right|\leq h_{0}e^{-\gamma_{0}r_{\ell k}}$
@@
Remark 2.1. We note that the constants $h_{0},\gamma_{0}$ in (2.7) may be chosen to be independent of atomic species. Additionally, (2.7) requires that the dependence of ${\mathcal{H}}(u)^{ab}_{\ell k}$ on site $m$ decays exponentially in $r_{\ell m}(u)+r_{mk}(u)$. That is, we are assuming long-range Coulomb interactions are screened, a common assumption in many practical tight binding models [10, 25, 32]. We assume this to separate Coulomb type long-range effects from those arising due to small band gaps.
@@ 2.3. Direct and Indirect Band Gaps
Next, we review the concepts of direct and indirect band gaps. For this purpose, we first recall the Bloch transform, which provides a decomposition of ${\mathcal{H}}^{\rm ref}$, denoted by ${\mathcal{H}}_{\boldsymbol{\xi}}$ for all $\bm{\xi}\in\Gamma^{*}$. For more details, see Appendix A. By applying an eigenvalue decomposition for ${\mathcal{H}}_{\boldsymbol{\xi}}$, we can obtain energy band structure of the system and subsequently distinguish between direct and indirect band gaps. The Bloch fibres of ${\mathcal{H}}^{\rm ref}$ are given by
(2.8) $\big[{\mathcal{H}}_{\boldsymbol{\xi}}\big]^{ab}_{\ell_{0}k_{0}}\coloneqq\sum_{\alpha\in\mathbb{Z}^{d}}[{\mathcal{H}}^{\mathrm{ref}}]^{ab}_{\ell_{0}+\mathsf{A}\alpha,k_{0}}e^{-i(\ell_{0}-k_{0}+\mathsf{A}\alpha)\cdot\bm{\xi}}\quad\forall\bm{\xi}\in\Gamma^{*},$
for all $\ell_{0},k_{0}\in\Lambda_{0}\coloneqq\Gamma\cap\Lambda$ and $1\leq a,b\leq N_{\mathrm{b}}$. The size of the matrix ${\mathcal{H}}_{\boldsymbol{\xi}}$ is $MN_{\mathrm{b}}\times MN_{\mathrm{b}}$. For $\ell\in\Lambda$, we write $\ell_{0}$ for the unique site in $\Lambda_{0}$ satisfying $\ell-\ell_{0}\in\mathsf{A}\mathbb{Z}^{d}$. Denoting the eigenvalues of ${\mathcal{H}}_{\boldsymbol{\xi}}$ by $\varepsilon_{1}(\bm{\xi})\leq\cdots\leq\varepsilon_{MN_{\mathrm{b}}}(\bm{\xi})$, we have
(2.9) $\sigma({\mathcal{H}}^{\mathrm{ref}})=\bigcup_{j=1}^{MN_{\mathrm{b}}}\left\{\varepsilon_{j}(\bm{\xi}):~\bm{\xi}\in\Gamma^{*}\right\},$
@@
We focus exclusively on systems with a positive band gap, such as insulators and semi-conductors, where the Fermi level, $\varepsilon_{\mathrm{F}}$, lies inside the band gap:
(Gap)   The Fermi level $\varepsilon_{\mathrm{F}}$ separates the energy bands into the valence bands and conduction bands [11]:
(2.10) $\varepsilon_{\mathrm{F}}\not\in\sigma({\mathcal{H}}^{\mathrm{ref}})\quad\text{and}\quad\varepsilon_{N_{0}}(\bm{\xi})<\varepsilon_{\mathrm{F}}<\varepsilon_{N_{0}+1}(\bm{\xi})\quad\forall\bm{\xi}\in\Gamma^{*},$
@@
In semiconductors, one can consider two types of band gaps $\mathsf{gap}_{-}\leq\mathsf{gap}_{+}$ defined as
(2.11) $\mathsf{gap}_{+}$ $\coloneqq\min_{\bm{\xi}\in\Gamma^{*}}\big(\varepsilon_{N_{0}+1}(\bm{\xi})-\varepsilon_{N_{0}}(\bm{\xi})\big),\qquad\text{and}$
(2.12) $\mathsf{gap}_{-}$ $\coloneqq\min_{\bm{\xi}\in\Gamma^{*}}\varepsilon_{N_{0}+1}(\bm{\xi})-\max_{\bm{\xi}\in\Gamma^{*}}\varepsilon_{N_{0}}(\bm{\xi}).$
Figure 1. Schematic of the direct and indirect band gaps.
Figure 2. The band structure of cubic $\rm{Mg_{2}Si}$ and cubic diamond carbon. In diamond, we have $\mathsf{gap}_{-}\approx\mathsf{gap}_{+}$, whereas in $\rm{Mg_{2}Si}$, we observe $\mathsf{gap}_{-}\ll\mathsf{gap}_{+}$.
@@ 3. Main Results
In this section, we present our main results on the nearsightedness property, which we sometimes call weak locality to distinguish it from locality of the mechanical response (strong locality).
At zero Fermi-temperature, the Fermi–Dirac distribution is given by $\chi_{(-\infty,\varepsilon_{\mathrm{F}})}$ and thus the density matrix reads
(3.1) $\rho(u)=\chi_{(-\infty,\varepsilon_{\mathrm{F}})}({\mathcal{H}}(u)),\quad u\in\mathrm{Adm}(\Lambda).$
For simplicity, we will denote $\rho^{\mathrm{ref}}:=\rho(\bm{0})$, and use $\left|\rho^{\mathrm{ref}}_{\ell k}\right|$ to refer to the Frobenius norm of each block for any $\ell,k\in\Lambda$.
We begin by considering the homogeneous lattice and present an improved estimate for the locality of the density matrix, explicitly capturing its linear dependence on $\mathsf{gap}_{+}$:
Theorem 3.1 (Nearsightedness). Suppose that $\Lambda$ has the form of (2.1) and ${\mathcal{H}}^{\mathrm{ref}}$ satisfies assumptions (TB) and (Gap). Then, there exist constants $C_{1},\eta_{+}>0$ such that
(3.2) $\left|\rho^{\mathrm{ref}}_{\ell k}\right|\leq\frac{C_{1}}{\mathsf{gap}_{+}}e^{-{\eta_{+}}r_{\ell k}},\quad\forall\ell,k\in\Lambda,$
where $\eta_{+}:=c_{1}\frac{\gamma_{0}}{h_{0}}\min\left\{h_{0},\gamma_{0}^{d}\,\mathsf{gap}_{+}\right\}$ and $c_{1},C_{1}>0$ only depend on $d,\mathfrak{m},M$ and $N_{\mathrm{b}}$.
@@
Remark 3.2 (Existing results). We briefly note here that (3.2) but with $\mathsf{gap}_{+}$ replaced with $\mathsf{gap}_{-}$ follows from a contour integral representation of the density matrix, together with Combes–Thomas type estimates on the resolvent [1, 7, 13, 15, 28]. That is, we let $\mathscr{C}$ be a simple, closed, positively oriented contour encircling $\sigma({\mathcal{H}}^{\mathrm{ref}})\cap(-\infty,\varepsilon_{\mathrm{F}})$ (depicted in Figure 3), write
(3.3) $\rho^{\mathrm{ref}}=\oint_{\mathscr{C}}\chi_{(-\infty,\varepsilon_{\mathrm{F}})}(z)\big(z-{\mathcal{H}}^{\mathrm{ref}}\big)^{-1}\frac{\mathrm{d}z}{2\pi i}=\oint_{\mathscr{C}}\big(z-{\mathcal{H}}^{\mathrm{ref}}\big)^{-1}\frac{\mathrm{d}z}{2\pi i},$
@@
(3.4) $\left|\big(z-{\mathcal{H}}^{\mathrm{ref}}\big)^{-1}_{\ell k}\right|\leq\frac{2}{\mathfrak{d}^{\rm ref}}e^{-\gamma_{\mathrm{CT}}(\mathfrak{d}^{\rm ref})r_{\ell k}}$
where $\gamma_{\mathrm{CT}}(\mathfrak{d}^{\rm ref})=c_{0}\frac{\gamma_{0}}{h_{0}}\min\{h_{0},\gamma_{0}^{d}\,\mathfrak{d}^{\rm ref}\}$ with $c_{0}=c_{0}(d,\mathfrak{m})$ depending only on the geometry. ∎
Figure 3. A schematic plot of the contour $\mathscr{C}$ encircling the valence bands denoted by the green strip, and the conduction bands denoted in blue.
Sketch of the proof of Theorem 3.
@@
$\rho^{\mathrm{ref}}_{\ell k}$ $=\oint_{\mathscr{C}}\bigg[\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{\star}}\left(z-{\mathcal{H}}_{\boldsymbol{\xi}}\right)^{-1}e^{i\bm{\xi}\cdot(\ell-k)}\mathrm{d}\bm{\xi}\bigg]_{\ell_{0}k_{0}}\frac{\mathrm{d}z}{2\pi i}$
(3.5) $=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\bigg[\oint_{\mathscr{C}+\nu(\bm{\xi})}\left(z-{\mathcal{H}}_{\boldsymbol{\xi}}\right)^{-1}\frac{\mathrm{d}z}{2\pi i}\bigg]_{\ell_{0}k_{0}}e^{i\bm{\xi}\cdot(\ell-k)}{\,{\rm d}}\bm{\xi}\eqqcolon\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\big[\widehat{\rho}(\bm{\xi})\big]_{\ell_{0}k_{0}}\,e^{i\bm{\xi}\cdot(\ell-k)}\mathrm{d}\bm{\xi},$
@@
where $\ell_{0},k_{0}\in\Lambda_{0}$ satisfy $\ell-\ell_{0},k-k_{0}\in\mathsf{A}\mathbb{Z}^{d}$ and $\nu:\Gamma^{*}\rightarrow\mathbb{C}$ is such that $\mathscr{C}+\nu(\bm{\xi})$ encircles $\{\varepsilon_{n}(\bm{\xi})\}_{n\leq N_{0}}$ and avoids $\{\varepsilon_{n}(\bm{\xi})\}_{n\geq N_{0}+1}$.
@@
We conclude (3.2) by showing that $\nu$ and thus $\widehat{\rho}$ extend to analytic functions on a strip of width proportional to $\mathsf{gap}_{+}$, and then apply a suitable Paley–Wiener estimate. Full details are provided in Section 6.1. ∎
@@
We now establish the stability of the nearsightedness estimates under lattice perturbations. Specifically, we extend our results to inhomogeneous systems characterized by finite-energy displacements. Since those perturbations must have finite energy-norm, they are necessarily localized, but with relatively slow (e.g., algebraic) decay [5].